OpenStudy (anonymous):

y = 4x - 2 y = x + 13

OpenStudy (eric_d):

Solve it simultaneously @nathanjames

OpenStudy (anonymous):

I assume you mean y=4x-2 and y=x+13 4x-2=x+13 4x-x=13+2 3x=15 x=5 therefore y=5+13.....y=18 ordered pair is (5,18)